# Nonlocal Contributions to the Turbulent Cascade in Magnetohydrodynamic Plasmas

### 19 Mar 2024 v2

@@ IV Conclusions
The exact law established by Politano and Pouquet [23, 24] assumes homogeneity, isotropy, and incompressibility - conditions that are fulfilled perhaps only locally in interplanetary space plasmas - representing to date the most robust statistical framework to characterize energy transfer and heating in the solar wind [7, 6]. Without diminishing its validity, our findings suggest that in some solar wind samples, the current analyses of observational data by the original P&P law and some of its surrogates might substantially overestimate the rate at which energy is dissipated (or further transferred across sub-proton scales) at the bottom of the MHD turbulent cascade developing in the solar wind. This leads to improved third-order moment law predictions characterizing the observed space plasma heating.
As mentioned in Sec. II, the additional nonlocal term arises due to the advection of the magnetic field by the velocity field and the corresponding transformation of the electric field ${\bf E}^{\prime}={\bf E}+\mu{\bf u}\times{\bf H}$. Hence, it would also be interesting to assess other transformations of the electric field, e.g., in the context of Hall MHD [62, 51]. Future work will be devoted to applying the law (18) and assessing the influence of nonlocality in solar wind measurements. Moreover, we aim at reconstructing velocity and magnetic fields from partial measurements by extending current velocity field reconstructions [65, 66] to MHD turbulence [67, 68, 69], exploiting observations of state-of-art ongoing [70, 71] and future multi-spacecraft solar wind mission [72, 73, 74, 75].
@@ V Acknowledgments
J.F. acknowledges fruitful discussions with H. Homann, A. Pumir, and J. Peinke. We gratefully acknowledge the Gauss Centre for Supercomputing (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS [76] at Jülich Supercomputing Centre (JSC). The simulation output was also analyzed on HPC facilities at the École Centrale de Lyon (PMCS2I) in Ecully (France) and the HPC Cluster STORM, located at the University of Oldenburg (Germany) and funded through the REACT-EU program by the NBank and the Ministry of Science and Culture (MWK) of the Lower Saxony State under grant number ZW7-85186744. R.M. acknowledges support from the project “EVENTFUL” (ANR-20-CE30-0011), funded by the French “Agence Nationale de la Recherche” - ANR through the program AAPG-2020.
@@ Appendix B Derivation of the Friedmann-Keller Hierarchy in Homogeneous and Isotropic MHD Turbulence
In this section, we recapitulate Chandrasekhar’s derivation [60] of the Friedmann-Keller hierarchy [77] for magnetic and velocity field correlation functions. In a statistical description of MHD turbulence, evolution equations for the two-point correlation tensors
@@
are obtained by the same procedure as in the hydrodynamic case [61] and were first derived by Chandrasekhar [60]. In the following, we are solely interested in evolution equations for the two-point velocity and magnetic field tensors (39) and (41). The evolution equation and invariant form of the cross helicity tensor (40), which is a skew-symmetric tensor, can be found in [60]. The procedure for the velocity correlation tensor (39) starts with multiplying Eq. (25) by $u_{j}^{\prime}=u_{j}({\bf x}^{\prime},t)$
@@
  • Pressure-velocity contributions vanish based on isotropy and mirror symmetry [78], which yields
@@
These two equations (and the additional equation for the cross helicity tensor of second order derived by Chandrasekhar [60]) are a generalization of the Friedmann-Keller hierarchy [77] to homogeneous and isotropic MHD turbulence.
@@
A further simplification of Eqs. (60) and (61) can be obtained from the invariant theory of homogeneous and isotropic turbulence [79]. Each tensor can be rewritten in terms of its longitudinal correlation function, which is a consequence of the incompressibility condition and is further derived in Appendices F,G, and H. The second-order tensors (46) and (50) both follow the tensorial form
@@
$\frac{\partial}{\partial t}Q_{kin}(r,t)$ $=$ $\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(r^{3}\frac{\partial}{\partial t}C_{r\;r}^{{\bf u}{\bf u}}(r,t),\right),\qquad\frac{\partial}{\partial t}Q_{mag}(r,t)=\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(r^{3}\frac{\partial}{\partial t}C_{r\;r}^{{\bf h}{\bf h}}(r,t),\right)$ (69)
$\frac{\partial}{\partial r_{k}}J_{k}^{kin}({\bf r},t)$ $=$ $-\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial}{\partial r}\left[r^{4}\left(C_{r\;r\;r}^{{\bf u}{\bf u}{\bf u}}-C_{r\;r\;r}^{{\bf h}{\bf h}{\bf u}}\right)\right]\right)\;,\qquad\frac{\partial}{\partial r_{k}}J_{k}^{mag}({\bf r},t)=-2\frac{1}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}(r,t)\right)\;,$ (70)
@@
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=$ $\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)+2\lambda\frac{1}{r^{4}}\frac{\partial}{\partial r}r^{4}\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\;.$ (73)
@@ Appendix D On the Equivalence of the Mixed Third-Order Correlations for Small Scale Separations
As first recognized by Chandrasekhar [60], the defining scalars of the mixed third-order correlation functions $C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)=\left\langle h_{i}({\bf x},t)h_{k}({\bf x},t)u_{j}({\bf x}^{\prime},t)\right\rangle$ and $A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)=\left\langle(h_{i}({\bf x},t)u_{k}({\bf x},t)-h_{k}({\bf x},t)u_{i}({\bf x},t))h_{j}({\bf x}^{\prime},t)\right\rangle$ are related to each other for small scale separations ${\bf r}$. This can be shown as follows: First, we consider the derivative of the symmetric tensor defined by Eq. (63)
@@
$=$ $-\left[\frac{C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{2r}+\frac{3}{4}\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right]_{r=0}\delta_{jk}$
$=$ $-\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right|_{r=0}\delta_{jk}\;.$ (74)
@@
$=$ $\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}\delta_{jk}\;.$ (75)
@@
$2\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}\delta_{jk}=-\left\langle h_{i}({\bf x},t)h_{j}({\bf x},t)\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right\rangle$
@@
$\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}=\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right|_{r=0}\;.$ (79)
@@ Appendix E Loitsiansky Invariants and their Implications for the Additional Source Term
In this section, we discuss further implications of Eqs. (8) and (9), namely the existence of certain invariants. As shown by Loitsiansky [80], in the hydrodynamic limit, the von Kármán-Howarth equation admits the invariant
@@
As shown by Chandrasekhar [60], this quantity should also be conserved in MHD turbulence, which can be seen by multiplying Eq. (8) by $r^{4}$ and subsequent integration
@@
Following Chandrasekhar [60], we introduce the vector potential ${\bf a}({\bf x},t)$ according to
@@ Appendix F Longitudinal and Transverse Correlation Functions
by $\frac{r_{i}}{r_{j}}$ and $\frac{r_{j}}{r}$. Assuming that $C_{i\,j}({\bf r},t)=\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\rangle$ is isotropic and mirror-symmetric [22] its general form is given by
@@
The bilinear form (92) can now be extended to a trilinear form [79]. A general isotropic and mirror-symmetric tensor of order three can be defined as
@@ G.1 Correlation Functions of Second Order
$\frac{\partial}{\partial r_{i}}C_{i\,j}({\bf r},t)=\left\langle\frac{\partial u_{i}({\bf x}+{\bf r},t)}{\partial r_{i}}u_{j}({\bf x},t),\right\rangle=0\;.$ (104)

### 10 Apr 2024 v3

@@ IV Conclusions
The exact law established by Politano and Pouquet Politano and Pouquet (1998a, b) assumes homogeneity, isotropy, and incompressibility - conditions that are fulfilled perhaps only locally in interplanetary space plasmas - representing to date the most robust statistical framework to characterize energy transfer and heating in the solar wind Marino and Sorriso-Valvo (2023); Marino et al. (2008). Without diminishing its validity, our findings suggest that in some solar wind samples, the current analyses of observational data by the original P&P law and some of its surrogates might substantially overestimate the rate at which energy is dissipated (or further transferred across sub-proton scales) at the bottom of the MHD turbulent cascade developing in the solar wind. This leads to improved third-order moment law predictions characterizing the observed space plasma heating.
As mentioned in Sec. II, the additional nonlocal term arises due to the advection of the magnetic field by the velocity field and the corresponding transformation of the electric field ${\bf E}^{\prime}={\bf E}+\mu{\bf u}\times{\bf H}$. Hence, it would also be interesting to assess other transformations of the electric field, e.g., in the context of Hall MHD Galtier (2008); Foldes et al. (2023). Future work will be devoted to applying the law (18) and assessing the influence of nonlocality in solar wind measurements. Moreover, we aim at reconstructing velocity and magnetic fields from partial measurements by extending current velocity field reconstructions Friedrich et al. (2022); Lübke et al. (2023) to MHD turbulence Friedrich (2020); Friedrich and Grauer (2020); Lübke et al. (2024), exploiting observations of state-of-art ongoing (Owen et al., 2020; Fox et al., 2016) and future multi-spacecraft solar wind mission (Klein et al., 2023; Pecora et al., 2023a, b; Retino et al., 2022).
@@ V Acknowledgments
J.F. acknowledges fruitful discussions with H. Homann, A. Pumir, and J. Peinke. We gratefully acknowledge the Gauss Centre for Supercomputing (www.gauss-centre.eu) for funding this project by providing computing time through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputer JUWELS Alvarez (2021) at Jülich Supercomputing Centre (JSC). The simulation output was also analyzed on HPC facilities at the École Centrale de Lyon (PMCS2I) in Ecully (France) and the HPC Cluster STORM, located at the University of Oldenburg (Germany) and funded through the REACT-EU program by the NBank and the Ministry of Science and Culture (MWK) of the Lower Saxony State under grant number ZW7-85186744. R.M. acknowledges support from the project “EVENTFUL” (ANR-20-CE30-0011), funded by the French “Agence Nationale de la Recherche” - ANR through the program AAPG-2020. M.W.  acknowledges support from the German Science Foundation DFG within the Collaborative Research Center SFB1491.
@@ Appendix B Derivation of the Friedmann-Keller Hierarchy in Homogeneous and Isotropic MHD Turbulence
In this section, we recapitulate Chandrasekhar’s derivation Chandrasekhar (1951) of the Friedmann-Keller hierarchy Keller and Friedman (1924) for magnetic and velocity field correlation functions. In a statistical description of MHD turbulence, evolution equations for the two-point correlation tensors
@@
are obtained by the same procedure as in the hydrodynamic case de Karman and Howarth (1938) and were first derived by Chandrasekhar Chandrasekhar (1951). In the following, we are solely interested in evolution equations for the two-point velocity and magnetic field tensors (39) and (41). The evolution equation and invariant form of the cross helicity tensor (40), which is a skew-symmetric tensor, can be found in Chandrasekhar (1951). The procedure for the velocity correlation tensor (39) starts with multiplying Eq. (25) by $u_{j}^{\prime}=u_{j}({\bf x}^{\prime},t)$
@@
  • Pressure-velocity contributions vanish based on isotropy and mirror symmetry Faust et al. (2015), which yields
@@
These two equations (and the additional equation for the cross helicity tensor of second order derived by Chandrasekhar Chandrasekhar (1951)) are a generalization of the Friedmann-Keller hierarchy Keller and Friedman (1924) to homogeneous and isotropic MHD turbulence.
@@
A further simplification of Eqs. (60) and (61) can be obtained from the invariant theory of homogeneous and isotropic turbulence Robertson (1940). Each tensor can be rewritten in terms of its longitudinal correlation function, which is a consequence of the incompressibility condition and is further derived in Appendices F,G, and H. The second-order tensors (46) and (50) both follow the tensorial form
@@
$\frac{\partial}{\partial t}Q_{kin}(r,t)$ $=$ $\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(r^{3}\frac{\partial}{\partial t}C_{r\;r}^{{\bf u}{\bf u}}(r,t)\right),\qquad\frac{\partial}{\partial t}Q_{mag}(r,t)=\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(r^{3}\frac{\partial}{\partial t}C_{r\;r}^{{\bf h}{\bf h}}(r,t)\right),$ (69)
$\frac{\partial}{\partial r_{k}}J_{k}^{kin}({\bf r},t)$ $=$ $-\frac{1}{2r^{2}}\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial}{\partial r}\left[r^{4}\left(C_{r\;r\;r}^{{\bf u}{\bf u}{\bf u}}-C_{r\;r\;r}^{{\bf h}{\bf h}{\bf u}}\right)\right]\right)\;,\qquad\frac{\partial}{\partial r_{k}}J_{k}^{mag}({\bf r},t)=\frac{2}{r^{2}}\frac{\partial}{\partial r}\left(r^{2}A_{[rt]t}^{\mathbf{u}\mathbf{h}{\mathbf{h}}}(r,t)\right)\;,$ (70)
@@
$\frac{\partial}{\partial t}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)=$ $-\frac{4}{r}A_{[rt]t}^{\mathbf{u}\mathbf{h}\mathbf{h}}(r,t)+2\lambda\frac{1}{r^{4}}\frac{\partial}{\partial r}r^{4}\frac{\partial}{\partial r}C_{r\;r}^{\mathbf{h}\mathbf{h}}(r,t)\;.$ (73)
@@ Appendix D On the Equivalence of the Mixed Third-Order Correlations for Small Scale Separations
As first recognized by Chandrasekhar Chandrasekhar (1951), the defining scalars of the mixed third-order correlation functions $C_{(ki)j}^{{\bf h}{\bf h}{\bf u}}({\bf r},t)=\left\langle h_{i}({\bf x},t)h_{k}({\bf x},t)u_{j}({\bf x}^{\prime},t)\right\rangle$ and $A_{[ki]j}^{{\bf u}{\bf h}{\bf h}}({\bf r},t)=\left\langle(h_{i}({\bf x},t)u_{k}({\bf x},t)-h_{k}({\bf x},t)u_{i}({\bf x},t))h_{j}({\bf x}^{\prime},t)\right\rangle$ are related to each other for small scale separations ${\bf r}$. This can be shown as follows: First, we consider the derivative of the symmetric tensor defined by Eq. (63)
@@
$=$ $\left[\frac{C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{r}+\frac{3}{2}\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right]_{r=0}\delta_{jk}$
$=$ $\frac{5}{2}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right|_{r=0}\delta_{jk}\;.$ (74)
@@
$=$ $\left.2\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}\delta_{jk}\;.$ (75)
@@
$4\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}\delta_{jk}=-\left\langle h_{i}({\bf x},t)h_{j}({\bf x},t)\frac{\partial u_{k}({\bf x},t)}{\partial x_{i}}\right\rangle$
@@
$\left.\frac{\partial A_{[rt]t}^{{\bf u}{\bf h}{\bf h}}(r,t)}{\partial r}\right|_{{r}=0}=-\frac{5}{4}\left.\frac{\partial C_{r\;r\;r}^{\mathbf{h}\mathbf{h}\mathbf{u}}(r,t)}{\partial r}\right|_{r=0}\;.$ (79)
@@ Appendix E Loitsiansky Invariants and their Implications for the Additional Source Term
In this section, we discuss further implications of Eqs. (8) and (9), namely the existence of certain invariants. As shown by Loitsiansky Loitsianskii (1939), in the hydrodynamic limit, the von Kármán-Howarth equation admits the invariant
@@
As shown by Chandrasekhar Chandrasekhar (1951), this quantity should also be conserved in MHD turbulence, which can be seen by multiplying Eq. (8) by $r^{4}$ and subsequent integration
@@
Following Chandrasekhar Chandrasekhar (1951), we introduce the vector potential ${\bf a}({\bf x},t)$ according to
@@ Appendix F Longitudinal and Transverse Correlation Functions
by $\frac{r_{i}}{r_{j}}$ and $\frac{r_{j}}{r}$. Assuming that $C_{i\,j}({\bf r},t)=\langle u_{i}({\bf x},t)u_{j}({\bf x}+{\bf r},t)\rangle$ is isotropic and mirror-symmetric Monin and Yaglom (2007) its general form is given by
@@
The bilinear form (92) can now be extended to a trilinear form Robertson (1940). A general isotropic and mirror-symmetric tensor of order three can be defined as
@@ G.1 Correlation Functions of Second Order
$\frac{\partial}{\partial r_{i}}C_{i\,j}({\bf r},t)=\left\langle\frac{\partial u_{i}({\bf x}+{\bf r},t)}{\partial r_{i}}u_{j}({\bf x},t)\right\rangle=0\;.$ (104)